On a see-saw mary, weight 600N balances john, weight 200N when she sits 1.5m away from the pivot.how far from the pivot is john?
1
Expert's answer
2020-11-03T10:28:20-0500

From the torques equality, we can write:


"600N\times 1.5m = 200N\times x"

where "x" is the distance from the pivot to John. Expressing "x", find:


"x = \dfrac{600N\times 1.5m}{200N} =4.5m"

Answer. 4.5 m.